Problem A 100 kg object that is at rest, pushes with a force of 250 Newtons, moving it in a straight line at a distance of 15 meters along a horizontal floor with a coefficient of kinetic friction equal to 0.25. Calculate:
a) the work of the applied force (WF)
b) the magnitude of the weight (w)
c) the magnitude of the kinetic friction force (Frk d) the work of the kinetic friction (WFrk)
e) the total work (Wtotal) f the final velocity of the object (v)
g) the final power of the applied force (Pfinal)
Answer #1

a)

wF = 250* 15 =3750 J

b)

e = 100* 9.8 = 980 N

c)

Frk = 0.25* 100* 9.8 = 245N

d)

wFrk = 245* 15 = - 3675 J

e)

W = 3750 - 3675 = 75 J

f)

0.5* 100* v^2 =75

v =1.2247 m/s

g)

Pf = 250 * 1.2247 =306.162 W

====

Comment before rate in case any doubt, will reply for sure.. goodluck